1) Units-of-activity Depreciation A truck acquired at a cost of $310,000 has an estimated residual value of $19,600, has an estimated useful life of 44,000 miles, and was driven 3,500 miles during the year. Determine the following. If required, round your answer for the depreciation rate to two decimal places. (a) The depreciable cost $ (b) The depreciation rate $ per mile (c) The units-of-activity depreciation for the year $

2)

Depreciation by Two Methods

A Kubota tractor acquired on January 8 at a cost of $297,000 has an estimated useful life of ten years. Assuming that it will have no residual value.

a. Determine the depreciation for each of the first two years by the straight-line method.

First Year Second Year
$ $

b. Determine the depreciation for each of the first two years by the double-declining-balance method. Do not round the double-declining balance rate. If required, round your final answer to the nearest dollar.

First Year Second Year
$ $
Answer #1

1)

a.Depreciable cost=(310,000-19600)=$290,400

b.depreciation rate=(290400/44000)=$6.6/mile

c.units-of-activity depreciation for the year=(6.6*3500)=$23,100